# Hyperbolic Random Graphs: Clique Number and Degeneracy with Implications for Colouring

### v1

@@ Appendix B Bounds on $\theta_{R}(\cdot,\cdot)$
Recall that $\theta_{R}(r,q)$ is the angle between points with radii $p$ and $q$ such that the distance between them is $R$. We apply Lemma B.1 to $\theta_{R}(r,q)$ to obtain the following.
Lemma B.3.
@@
Since $2e^{R-r-q}\in\left(0,2e^{-\Delta}\right]$ we can apply Lemma B.1. Evaluating the $\arccos$ for $r+q=R+\Delta$ for the domains of our interest we obtain
@@
Note that, by our hypothesis that $q\leq r\leq R/2+\mathcal{O}(1)$, there exists an $\varepsilon>0$, such that $2e^{R-x-y}+e^{-R-x-y}-\left(e^{-2x}+e^{-2r}\right)\in\left[\varepsilon,2\right)$. Thus, we can and we will apply Lemma B.1. We note that, for all $\varepsilon>0$ there exists a constant $\delta^{\prime}>0$ such that $\arccos(1-\varepsilon)\sqrt{\frac{2}{\varepsilon}}=2+\delta^{\prime}$. Hence there exists another constant $\delta$ such that $0<\delta<\delta^{\prime}$, for which we get
@@
We also consider vertices that have a sub-constant distance to the core. The statement here tells us that when $r$ approaches $R/2$ from above, the angle $\theta_{R}(r,R/2)$ approaches $\pi$ from below.
Lemma B.5.
Proof B.6.
@@
$\theta_{R}(r,x)$ $\geq\pi-2\sqrt{\log^{k}(n)n^{\alpha-1}},$
@@
Appendix C Lemma 3.4
Lemma B.3 allows us to bound the probability of a vertex $v$ belonging to the inner-neighbourhood area of a vertex $u$. We remark that these bounds are sharper than the ones obtained in [7, Lemma 3] for points with radius $R/2+\Theta(1)$. In contrast, for points with radius $R/2+\omega(1)$ the bound in [7, Lemma 3] is superior. Moreover, notice that for vertices with radius $r\leq R/2$, the inner-neighbourhood is simply $V\cap\mathcal{B}_{0}(r)$.
Lemma C.1.
@@
By Lemma B.3 we have $\theta_{R}(r,q)\leq\pi\sqrt{e^{(R-r-x)}}$ for $x+r\geq R$. Thus, for every $x\in[R-r,r]$, there exists a $\tau\leq\pi$ such that
@@
Notice that we can split the integral of the domain $[R-r,r]$ into further sub domains such that for domain $[s,t]$ we get $s+r\in R+\Theta(1)$. This allows to apply different bounds for $\theta_{R}(s,r)$ as shown in Lemma B.3, giving a more refined upper bounds on $\tau$. Thus, we apply Lemma B.3 to obtain
@@
$\quad+\operatorname{\mathbf{1}}_{\{r\geq R/2\}}\frac{\alpha e^{(R-r)/2}}{(\cosh(\alpha R)-1)}\int_{R-r}^{\min(r,R-r+\log(4/3))}e^{-x/2}\sinh(\alpha x)\,\mathrm{d}x$ (5)
$\quad+\operatorname{\mathbf{1}}_{\{r\geq R/2+\log(\sqrt{4/3})\}}\frac{{4}\alpha e^{(R-r)/2}}{{3\sqrt{3}}(\cosh(\alpha R)-1)}\int_{R-r+\log(4/3)}^{\min(r,{R-r+\log(2)})}e^{-x/2}\sinh(\alpha x)\,\mathrm{d}x$ (6)
$\quad+\operatorname{\mathbf{1}}_{\{r\geq R/2+\log(\sqrt{2})\}}\frac{\alpha e^{(R-r)/2}}{\sqrt{2}(\cosh(\alpha R)-1)}\int_{R-r+\log(2)}^{{\min(r,R-r+2\log(2))}}e^{-x/2}\sinh(\alpha x)\,\mathrm{d}x$ (7)
$\quad+\operatorname{\mathbf{1}}_{\{r\geq R/2+2\log(\sqrt{2})\}}\frac{2\alpha e^{(R-r)/2}}{3(\cosh(\alpha R)-1)}\int_{R-r+2\log(2)}^{r}e^{-x/2}\sinh(\alpha x)\,\mathrm{d}x.$ (8)
@@
We proceed by taking care of each line of Equations 4, 5, 6, 7 and 8 separately. We start by considering Equation 5. Recall that $\tau\leq\pi$ for $r+s\geq R$. Using (10) and (9) for the cases $\Delta<\log(\sqrt{4/3})$ and $\Delta\geq\log(\sqrt{4/3})$ respectively and using that for $r\leq R/2$ we get $\operatorname{\mathbf{1}}_{\{r\geq R/2\}}=0$ gives us
Equation 5$\leq\left(1+\Theta\left(e^{-\alpha R}\right)\right)\frac{\alpha e^{-\alpha r}}{(\alpha-1/2)}\cdot\begin{cases}0&\text{for }r\leq R/2,\\ {\left(e^{\frac{1}{2}(2\alpha-1)(2r-R)}-1\right)}&\text{for }0<\Delta<\log(\sqrt{4/3}),\\ \left(\left(\frac{4}{3}\right)^{(\alpha-1/2)}-1\right)&\text{for }\Delta\geq\log(\sqrt{4/3}).\end{cases}$
Moving on to Equations 6 and 7 , we get in similar fashion
@@
$\frac{4\alpha e^{-\alpha r}}{3\sqrt{3}(\alpha-1/2)}\cdot\begin{cases}0&\text{for }\Delta\leq\log(\sqrt{4/3}),\\ {\left(e^{\frac{1}{2}(2\alpha-1)(2r-R)}-\left(\frac{4}{3}\right)^{(\alpha-1/2)}\right)}&\text{for }\log(\sqrt{4/3})<\Delta<\log(\sqrt{2}),\\ \left(2^{(\alpha-1/2)}-\left(\frac{4}{3}\right)^{(\alpha-1/2)}\right)&\text{for }\Delta\geq\log(\sqrt{2}),\end{cases}$
@@
$\frac{\alpha e^{-\alpha r}}{\sqrt{2}(\alpha-1/2)}\cdot\begin{cases}0&\text{for }\Delta\leq\log(\sqrt{2}),\\ {\left(e^{\frac{1}{2}(2\alpha-1)(2r-R)}-2^{(\alpha-1/2)}\right)}&\text{for }\log(\sqrt{2})<\Delta<2\log(\sqrt{2}),\\ \left(2^{(2\alpha-1)}-2^{(\alpha-1/2)}\right)&\text{for }\Delta\geq 2\log(\sqrt{2}).\end{cases}$
@@
$\frac{2\alpha e^{-\alpha r}}{3(\alpha-1/2)}\cdot\begin{cases}0&\text{for }\Delta\leq 2\log(\sqrt{2}),\\ {\left(e^{\frac{1}{2}(2\alpha-1)(2r-R)}-2^{(2\alpha-1)}\right)}&\text{else.}\end{cases}$
@@
Proof C.4.
The calculations are similar but simpler to those for Lemma C.1. Recall that
@@
This, in conjunction with Lemma B.3 yields for a constant $\delta>0$
@@
Appendix D Theorem 3.5
The following statement provides the radius $r^{*}$ that maximises (expected) inner-degree.
Lemma D.1.
Let $r^{*}$ be the radial coordinate of the point in $\mathcal{D}_{R}$ that maximizes the measure of the inner-ball. Then $r^{*}=R/2+\log\left(\frac{\alpha\eta}{\gamma(1-\alpha)}\right)/(2\alpha-1)$.
Proof D.2.
Using the expression of $\mu(\mathcal{I}(r))$ in Lemma 3.4,
See 3.5
Proof D.3.
We shall apply Section 3 to upper bound the degeneracy by the maximal inner-degree. We accomplish this via the neighbourhood of a vertex that is formed within the inner-ball. Recall that Lemma D.1 gives us the radius $r^{*}$ by which the inner-neighbourhood $\mathcal{I}(r)$ is maximal. Moreover, by Lemma C.1 we have an upper-bound on the inner-neighbourhood. Thus we upper bound the measure of any inner-ball by plugging $r^{*}$ into the upper bound of $\mu(\mathcal{I}(r))$ to see that
@@
Case 1 [$101/2<\alpha<9/$ 10]: Recall that by Lemma D.1, $r^{*}$ is monotonically increasing in $\alpha$. Then, letting $\alpha$ approach $1/2$ from above, and setting $\gamma=\frac{1}{\sqrt{2}}$ and $\eta=\frac{1}{2\alpha}-(1-\frac{4}{3\sqrt{3}})(\frac{4}{3})^{(\alpha-1/2)}-(\frac{4}{3\sqrt{3}}-\frac{1}{\sqrt{2}})2^{(\alpha-1/2)}$, we get $r^{*}\geq R/2+1/2>R/2+\frac{1}{2}\log(2).$ This satisfies all necessary conditions for Lemma C.1 when $\Delta\geq\log(\sqrt{2})$. Applying (11) and noticing that it is monotonically increasing in $\alpha$,
@@
Case 2 [9/$19/10\leq\alpha<$ 1]: Taking $\gamma=\frac{2}{3}$ and $\eta=\frac{1}{2\alpha}-(1-\frac{4}{3\sqrt{3}})(\frac{4}{3})^{(\alpha-1/2)}-(\frac{4}{3\sqrt{3}}-\frac{1}{\sqrt{2}})2^{(\alpha-1/2)}-(\frac{1}{\sqrt{2}}-\frac{2}{3})2^{(2\alpha-1)}$, we have via Lemma D.1 that for $\alpha\geq 9/10$,
@@
Then, for such $\gamma$, $\eta$ and $\alpha$ we again obtain (12), but now for 9/$19/10\leq\alpha<$ 1. The proof for the upper bound is then finished by another combination of a Chernoff and union bound, similarly to the previous case.
@@
Taking $\gamma=\frac{2}{\pi}$ and $\eta=\frac{2}{\pi}-\frac{(\alpha-1/2)}{\alpha}$ from Lemma C.3 and plugging into Lemma D.1, we get that
@@
Making use of Lemma C.3 in conjunction with $r^{*}=R/2+\frac{\log\left(\frac{4\alpha+\pi-2\pi\alpha}{4(1-\alpha)}\right)}{2\alpha-1}$ and $\xi=\log\left(1+\frac{\log^{2}(n)}{n^{1-\alpha}}\right)/\alpha$ we have
@@
Appendix E Theorem 3.8
See 3.8
Proof E.1.
@@
We construct a colouring using the degeneracy. We do so following the lines drawn in [29] and compute a smallest-last vertex ordering. To bound the run time of the algorithm we apply [29, Lemma 1] that states a run time of $\mathcal{O}(|E|+|V|)$. The average degree for hyperbolic random graph is $\mathcal{O}(1)$ w.e.h.p. (see [18, Theorem 2.3]). Hence $|E|\in\mathcal{O}(n)$ w.e.h.p. and by definition $|V|=n$. We conclude that it requires at most $\mathcal{O}(n)$ time w.e.h.p. to compute the smallest-last vertex ordering of a hyperbolic random graph.
@@
Having the smallest-last vertex ordering in hand we colour the graph in the following fashion: Let $v_{1},\cdots,v_{n}$ be the smallest-last vertex ordering of the vertices. Iterating through the vertices in this order, let $i\in[n]$ and colour vertex $v_{i}$ at time step $i$ with the smallest colour available that previously has not been assigned to any neighbour of $v_{i}$. More formally, consider the induced subgraph $G[V_{i}]$ where $V_{i}=\{v_{1},v_{2},..v_{i}\}$ and set the colour for $v_{i}$ to be the colour $f(v_{i})=\min{(\{\mathbb{N}\setminus N_{i}\})}$ where $N_{i}=\{f(u)|u\in V_{i}\wedge\{u,v_{i}\}\in E(G)\}$. This concludes the description of the algorithm.
@@
The correctness of the algorithm is immediate by observing that when we assign a colour to a vertex $v$, this colour is different from the colours of all previously coloured neighbours of $v$. Notice that, given the smallest-last vertex ordering $v_{1},\cdots,v_{n}$, the colouring can be constructed in $\mathcal{O}(n)$ w.e.h.p. To back up this claim, we recall that $|E|\in\mathcal{O}(n)$ w.e.h.p. Then iterating through $n$ vertices, we need to check all neighbours. Thus, for each edge $\{u,v\}$, the algorithm checks the colour assigned to $u$ and assigned $v$ when $v=v_{i}$ and $u=v_{i}$ respectively. Again, we conclude with a run time of $\mathcal{O}(|E|+|V|)$ which boils down to $\mathcal{O}(n)$ w.e.h.p. This concludes the correctness and our claimed running time $\mathcal{O}(n)$.
@@
We finish the proof by showing the claimed approximation ratio of $4/3$. To this end, we first consider the number of colours used by our algorithm. Recall that, by Theorem 3.5, the degeneracy of a hyperbolic random graph to be at most $\kappa(G)\leq(1+o(1))\left(\frac{4}{3}\right)^{\alpha}\sigma(G)$ w.e.h.p. By [29], the number of colours used via the smallest-last vertex ordering is upper bounded by the degeneracy $\kappa(G)$. Moreover, the minimal amount of colours for $G$ are given by its chromatic number $\chi(G)$. Thus approximation ratio is given by $\frac{\kappa(G)}{\chi(G)}\leq\frac{\kappa(G)}{\omega(G)}$ since $\chi(G)\geq\omega(G)$. A look at Section 3 reminds us that $\omega(G)\geq\sigma(G)$. Since the upper bound on $\kappa(G)$ and the lower bound on $\omega(G)$ holds w.e.h.p., a union over their complementary events does not occur w.e.h.p. The $\sigma(G)$ for $\kappa(G)/\omega(G)$ cancels out yielding the desired approximation factor.
Appendix F Theorem 4.10
Theorem F.1.
$D\geq 2\sinh^{-1}\left(\sqrt{\frac{d+1}{2d}}\sinh(r)\right).$
See 4.9
Proof F.2.
Using that $d=2$, we directly get from Theorem F.1 for diameter $R$ that $\frac{R}{2}\geq\sinh^{-1}(\sqrt{3/4}\sinh(r)).$ Rearranging and using for $x\in\mathbb{R}$ that $\sinh(x)=\frac{1}{2}e^{x}(1-e^{-2x})$ yields
Appendix G Theorem 5.3

### v2

@@ A.1 Bounds on $\theta_{R}(\cdot,\cdot)$
Recall that $\theta_{R}(r,q)$ is the angle between points with radii $p$ and $q$ such that the distance between them is $R$. We apply Lemma A.1 to $\theta_{R}(r,q)$ to obtain the following.
Lemma A.3.
@@
Since $2e^{R-r-q}\in\left(0,2e^{-\Delta}\right]$ we can apply Lemma A.1. Evaluating the $\arccos$ for $r+q=R+\Delta$ for the domains of our interest we obtain
@@
Note that, by our hypothesis that $q\leq r\leq R/2+\mathcal{O}(1)$, there exists an $\varepsilon>0$, such that $2e^{R-x-y}+e^{-R-x-y}-\left(e^{-2x}+e^{-2r}\right)\in\left[\varepsilon,2\right)$. Thus, we can and we will apply Lemma A.1. We note that, for all $\varepsilon>0$ there exists a constant $\delta^{\prime}>0$ such that $\arccos(1-\varepsilon)\sqrt{\frac{2}{\varepsilon}}=2+\delta^{\prime}$. Hence there exists another constant $\delta$ such that $0<\delta<\delta^{\prime}$, for which we get
@@
We also consider vertices that have a sub-constant distance to the core in Proposition 4.8.
See 4.7
Proof A.5.
@@
$\theta_{R}(r,x)$ $\geq\pi-2\sqrt{\log^{k}(n)n^{\alpha-1}}.$
@@
A.2 Bounds on the inner-neighbourhood
Lemma A.3 allows us to bound the probability of a vertex $v$ belonging to the inner-neighbourhood area of a vertex $u$. We remark that these bounds are sharper than the ones obtained in [7, Lemma 3] for points with radius $R/2+\Theta(1)$. In contrast, for points with radius $R/2+\omega(1)$ the bound in [7, Lemma 3] is superior.
Lemma A.6.
@@ A.2 Bounds on the inner-neighbourhood
By Lemma A.3 we have $\theta_{R}(r,q)\leq\pi\sqrt{e^{(R-r-x)}}$ for $x+r\geq R$. Thus, for every $x\in[R-r,r]$, there exists a $\tau\leq\pi$ such that
@@
Notice that we can split the integral of the domain $[R-r,r]$ into further sub domains such that for domain $[s,t]$ we get $s+r\in R+\Theta(1)$. This allows to apply different bounds for $\theta_{R}(s,r)$ as shown in Lemma A.3, giving a more refined upper bounds on $\tau$. Thus, we apply Lemma A.3 to obtain
@@
$\quad+\operatorname{\mathds{1}}_{\{r\geq R/2\}}\frac{\alpha e^{(R-r)/2}}{(\cosh(\alpha R)-1)}\int_{R-r}^{\min(r,R-r+\log(4/3))}e^{-x/2}\sinh(\alpha x)\,\mathrm{d}x$ (6)
$\quad+\operatorname{\mathds{1}}_{\{r\geq R/2+\log(\sqrt{4/3})\}}\frac{{4}\alpha e^{(R-r)/2}}{{3\sqrt{3}}(\cosh(\alpha R)-1)}\int_{R-r+\log(4/3)}^{\min(r,{R-r+\log(2)})}e^{-x/2}\sinh(\alpha x)\,\mathrm{d}x$ (7)
$\quad+\operatorname{\mathds{1}}_{\{r\geq R/2+\log(\sqrt{2})\}}\frac{\alpha e^{(R-r)/2}}{\sqrt{2}(\cosh(\alpha R)-1)}\int_{R-r+\log(2)}^{{\min(r,R-r+2\log(2))}}e^{-x/2}\sinh(\alpha x)\,\mathrm{d}x$ (8)
$\quad+\operatorname{\mathds{1}}_{\{r\geq R/2+2\log(\sqrt{2})\}}\frac{2\alpha e^{(R-r)/2}}{3(\cosh(\alpha R)-1)}\int_{R-r+2\log(2)}^{r}e^{-x/2}\sinh(\alpha x)\,\mathrm{d}x.$ (9)
@@
We proceed by taking care of each line of Equations 5, 6, 7, 8 and 9 separately. We start by considering Equation 6. Recall that $\tau\leq\pi$ for $r+s\geq R$. Using (11) and (10) for the cases $\Delta<\log(\sqrt{4/3})$ and $\Delta\geq\log(\sqrt{4/3})$ respectively and using that for $r\leq R/2$ we get $\operatorname{\mathds{1}}_{\{r\geq R/2\}}=0$ gives us
Equation 6$\leq\left(1+\Theta\left(e^{-\alpha R}\right)\right)\frac{\alpha e^{-\alpha r}}{(\alpha-1/2)}\begin{cases}0&\text{for }r\leq R/2,\\ {\left(e^{\frac{1}{2}(2\alpha-1)(2r-R)}-1\right)}&\text{for }0<\Delta<\log(\sqrt{4/3}),\\ \left(\left(\frac{4}{3}\right)^{(\alpha-1/2)}-1\right)&\text{for }\Delta\geq\log(\sqrt{4/3}).\end{cases}$
Moving on to Equations 7 and 8 , we get in similar fashion
@@
$\frac{4\alpha e^{-\alpha r}}{3\sqrt{3}(\alpha-1/2)}\begin{cases}0&\text{for }\Delta\leq\log(\sqrt{4/3}),\\ {\left(e^{\frac{1}{2}(2\alpha-1)(2r-R)}-\left(\frac{4}{3}\right)^{(\alpha-1/2)}\right)}&\text{for }\log(\sqrt{4/3})<\Delta<\log(\sqrt{2}),\\ \left(2^{(\alpha-1/2)}-\left(\frac{4}{3}\right)^{(\alpha-1/2)}\right)&\text{for }\Delta\geq\log(\sqrt{2}),\end{cases}$
@@
$\frac{\alpha e^{-\alpha r}}{\sqrt{2}(\alpha-1/2)}\begin{cases}0&\text{for }\Delta\leq\log(\sqrt{2}),\\ {\left(e^{\frac{1}{2}(2\alpha-1)(2r-R)}-2^{(\alpha-1/2)}\right)}&\text{for }\log(\sqrt{2})<\Delta<2\log(\sqrt{2}),\\ \left(2^{(2\alpha-1)}-2^{(\alpha-1/2)}\right)&\text{for }\Delta\geq 2\log(\sqrt{2}).\end{cases}$
@@
$\frac{2\alpha e^{-\alpha r}}{3(\alpha-1/2)}\begin{cases}0&\text{for }\Delta\leq 2\log(\sqrt{2}),\\ {\left(e^{\frac{1}{2}(2\alpha-1)(2r-R)}-2^{(2\alpha-1)}\right)}&\text{else.}\end{cases}$
@@
Proof A.9.
The calculations are similar but simpler to those for Lemma A.6. Recall that
@@
This, in conjunction with Lemma A.3 yields for a constant $\delta>0$
@@
A.3 Bounds on the degeneracy
See 3.7
Proof A.10.
We shall apply Section 3 to upper bound the degeneracy by the maximal inner-degree. We accomplish this via the neighbourhood of a vertex that is formed within the inner-ball. Recall that Lemma 3.5 gives us the radius $r^{*}$ by which the inner-neighbourhood $\mathcal{I}(r)$ is maximal. Moreover, by Lemma A.6 we have an upper-bound on the inner-neighbourhood. Thus we upper bound the measure of any inner-ball by plugging $r^{*}$ into the upper bound of $\mu(\mathcal{I}(r))$ to see that
@@ A.3 Bounds on the degeneracy
Case 1 [$101/2<\alpha<9/$ 10]: Recall that by Lemma 3.5, $r^{*}$ is monotonically increasing in $\alpha$. Then, letting $\alpha$ approach $1/2$ from above, and setting $\gamma=\frac{1}{\sqrt{2}}$ and $\eta=\frac{1}{2\alpha}-(1-\frac{4}{3\sqrt{3}})(\frac{4}{3})^{(\alpha-1/2)}-(\frac{4}{3\sqrt{3}}-\frac{1}{\sqrt{2}})2^{(\alpha-1/2)}$, we get $r^{*}\geq R/2+1/2>R/2+\frac{1}{2}\log(2).$ This satisfies all necessary conditions for Lemma A.6 when $\Delta\geq\log(\sqrt{2})$. Applying (12) and noticing that it is monotonically increasing in $\alpha$,
@@
Case 2 [9/$19/10\leq\alpha<$ 1]: Taking $\gamma=\frac{2}{3}$ and $\eta=\frac{1}{2\alpha}-(1-\frac{4}{3\sqrt{3}})(\frac{4}{3})^{(\alpha-1/2)}-(\frac{4}{3\sqrt{3}}-\frac{1}{\sqrt{2}})2^{(\alpha-1/2)}-(\frac{1}{\sqrt{2}}-\frac{2}{3})2^{(2\alpha-1)}$, we have via Lemma 3.5 that for $\alpha\geq 9/10$,
@@
Then, for such $\gamma$, $\eta$ and $\alpha$ we again obtain (13), but now for 9/$19/10\leq\alpha<$ 1.
The proof for the upper bound is then finished by another combination of a Chernoff and union bound, similarly to the previous case.
@@
Taking $\gamma=\frac{2}{\pi}$ and $\eta=\frac{2}{\pi}-\frac{(\alpha-1/2)}{\alpha}$ from Lemma A.8 and plugging into Lemma 3.5, we get that
@@
Making use of Lemma A.8 in conjunction with $r^{*}=R/2+\frac{\log\left(\frac{4\alpha+\pi-2\pi\alpha}{4(1-\alpha)}\right)}{2\alpha-1}$ and $\xi=\log\left(1+\frac{\log^{2}(n)}{n^{1-\alpha}}\right)/\alpha$ we have
@@
A.4 Approximation algorithm
See 3.10
Proof A.11.
@@ A.4 Approximation algorithm
We construct a colouring using the degeneracy. We do so following Matula and Beck [30] and compute a smallest-last vertex ordering. To bound the run time of the algorithm we apply [30, Lemma 1] that states a run time of $\mathcal{O}(|E|+|V|)$. The average degree for hyperbolic random graph is $\mathcal{O}(1)$ w.e.h.p. (see [24, Corollary 17]). Hence $|E|\in\mathcal{O}(n)$ w.e.h.p. and by definition $|V|=n$. We conclude that it requires at most $\mathcal{O}(n)$ time w.e.h.p. to compute the smallest-last vertex ordering of a hyperbolic random graph.
@@
Having the smallest-last vertex ordering in hand we colour the graph in the following greedy fashion: let $v_{1},\cdots,v_{n}$ be the smallest-last vertex ordering of the vertices. Iterating through the vertices in this order, let $i\in[n]$ and colour vertex $v_{i}$ at time step $i$ with the smallest colour available that previously has not been assigned to any neighbour of $v_{i}$. More formally, consider the induced subgraph $G[V_{i}]$ where $V_{i}=\{v_{1},v_{2},..v_{i}\}$ and set the colour for $v_{i}$ to be the colour $f(v_{i})=\min{(\{\mathbb{N}\setminus N_{i}\})}$ where $N_{i}=\{f(u)|u\in V_{i-1}\wedge\{u,v_{i}\}\in E(G)\}$. This concludes the description of the algorithm.
@@
The correctness of the algorithm is immediate by observing that when we assign a colour to a vertex $v_{i}$, this colour is different from the colours of all previously coloured neighbours of $v_{i}$. Notice that, given the smallest-last vertex ordering $v_{1},\cdots,v_{n}$, the colouring can be constructed in $\mathcal{O}(n)$ w.e.h.p. To back up this claim, we recall that $|E|\in\mathcal{O}(n)$ w.e.h.p. and that the algorithm iterates through $n$ vertices while it checks the colour of all neighbours. Thus, for each edge $\{u,v_{i}\}\in E$, the algorithm checks the colour assigned to $u$ at step $i$ and over all iterations these are $|E|\in\mathcal{O}(n)$ checks w.e.h.p. Again, we conclude with a run time of $\mathcal{O}(|E|+|V|)$ which boils down to $\mathcal{O}(n)$ w.e.h.p. This concludes the correctness and our claimed running time $\mathcal{O}(n)$.
@@
We finish the proof by showing the claimed approximation ratio of $(4/3)^{\alpha}$. To this end, we first consider the number of colours used by our algorithm. Recall that, by Theorem 3.7, the degeneracy of a hyperbolic random graph is at most $\kappa(G)\leq(1+o(1))\left(\frac{4}{3}\right)^{\alpha}\sigma(G)$ w.e.h.p. By [30, Lemma 4], the number of colours used via the smallest-last vertex ordering is upper bounded by the degeneracy $\kappa(G)+1$. Moreover, the minimal amount of colours for $G$ are given by its chromatic number $\chi(G)$. Thus, an approximation ratio is given by $\frac{\kappa(G)+1}{\chi(G)}\leq\frac{\kappa(G)+1}{\sigma(G)}$ since $\chi(G)\geq\sigma(G)$ by Section 2. Since the upper bound on $\kappa(G)$ and the lower bound on $\omega(G)$ holds w.e.h.p., a union over their complementary events does not occur w.e.h.p. The $\sigma(G)$ for $\kappa(G)/\sigma(G)$ cancels out yielding the desired approximation factor w.e.h.p.
A.5 Geometric inhomogeneous random graphs